OpenStudy (anonymous):

A,B,C are Im of z=-1+5i, w=1-i, u=2k-1-(5k-3)i , k ε R . a)Find the vectors AB and BC. b)for which k A,B,C are on the same lane.

OpenStudy (anonymous):

a. AB=(1-i)-(-1+5i)=2-6i BC=2K-1-(5K-3)i-(1-i)=2k-2-(5k-4)

OpenStudy (anonymous):

It's not going like this: A(-1,5) , B(1,-1) and C(2k-1,-5k+3) ?